# Tracing Vulnerability Propagation Across Open Source Software Ecosystems

### v1

@@ 1 Introduction
The paper continues and advances the empirical vulnerability coordination research theme [7, 10] in a software ecosystem context; the interest is to better understand how vulnerabilities propagate across ecosystems. The ecosystem context places the paper into a large branch that has examined the software and cyber security of particularly programming language ecosystems, including with respect to a recent surge of malware uploads to these ecosystems [11]. With respect to software traceability research, which too is a large branch, the paper operates in a post-release traceability context because all software observed has already been released [4, 9]. Patching and otherwise handling of vulnerabilities is a classical example of typical post-release software engineering activities [6, 15], and, hence, it is not a surprise that also vulnerabilities have been traced [1, 14].
The paper’s practical relevance can be motivated by the Cyber Resilience Act (CRA) recently agreed upon in the European Union.111 Regulation (EU) 2024/2847. Among the regulation’s essential cyber security requirements are legal obligations to follow a coordinated vulnerability disclosure policy, supply security patches, and to only ship products without known vulnerabilities. As open source software is widely used also in the commercial software industry, fulfilling the obligations requires tracking, coordinating, and tracing of vulnerabilities also with respect to open source software components, including those distributed through software ecosystems.
@@ 2.1 Data
A note should be also made about the ecosystem term; it refers to the corresponding concept used in the OSV database.333 https://google.github.io/osv.dev/data/#covered-ecosystems These ecosystems cover programming language ecosystems, such as PyPI for Python or Packagist for PHP, Linux distributions, such as Red Hat and Ubuntu, software testing frameworks, such as OSS-Fuzz, and large hosting services, such as GitHub for which GitHub Actions were merged with the main GitHub ecosystem entries. This terminological choice can be justified on the grounds that there is no consensus in the literature about definitions [3]. The choice also maintains coherence with the OSV.
@@ 2.2 Methods
[…]
@@ 3 Results
The propagation sequences elaborated in (3) and (4) provide a good way to start the dissemination of the empirical results. Thus, the top-30 sequences (in terms of frequency) are shown in Fig. 2. (For presentation purposes, the figure shows also cases with $|\sigma_{c}|=1$.) In total, a little over four thousand unique sequences were identified for the about $85$ thousand unique CVEs. This amount alone testifies that the propagation can be quite complex among the $28$ ecosystems sampled. Both the mean and median are six ecosystems per a CVE on average. However, GitHub alone leads the ranking in Fig. 2; nearly $14\%$ of the CVEs observed were reported only in GitHub and nowhere else. GitHub is also present in many traces with a $|\sigma_{c}|=2$, meaning that many other ecosystems, including both programming language ecosystems and Linux distributions, have tended to pick CVE-referenced vulnerabilities that were initially reported on GitHub.
@@
The second place in the ranking is taken by CVEs first reported in Debian and then in Ubuntu. This observation is logical because most of Ubuntu’s packages are distributed also in Debian, and largely also maintained therein. Having said that, there are also quite a few CVEs that were reported either in Debian or Ubuntu alone. Although further validation would be required, this observation could be taken to support an argument that there might be still room for better coordination and synchronization. The same point extends toward other Linux distributions who too seem to often pick CVEs that were first reported in Debian. In this regard, it also worth emphasizing that Debian, Ubuntu, Red Hat, and SUSE, among a few others, such as the Python Software Foundation who maintains PyPI, are official CVE numbering authorities (CNAs) unlike Alpine Linux, Mageia, Rocky Linux, and some others [8]. Because the dataset’s scope was restricted to CVE-referenced vulnerabilities, and because CVEs require coordination of their own [10], the propagation sequences observed may be partially—but unlikely fully—explained by the CNAs present in the sequences.
@@
Regarding $\textmd{H}_{2}$, the Bonferroni-corrected and variance-adjusted $t$-tests indicate that for ten ecosystems out of $27$ the differences between the means are not different from zero, meaning that a presence or an absence of a given $a$ does not affect the averages of the given traceability delays.444 CRAN was excluded because only a single CVE has been reported for it. Among these ten ecosystems are GitHub, Maven, openSUSE, Packagist, PyPI, RubyGems, and Ubuntu. In contrast, statistically significant differences are present for Debian, Mageia, Red Hat, and SUSE for which their presences indicate longer delays. If the propagation sequences have involved Go or npm, the delays have been shorter. In general, these test results indicate that there is no clear-cut pattern among the ecosystems with respect to the traceability delays, including regarding the CNAs.
@@ 4 Conclusion
Regarding the results, (1) the propagation has generally been rather complex, as demonstrated by over four thousand unique propagation sequences for the $28$ ecosystems and about $85$ thousand unique CVEs observed. In terms of a frequency ranking, however, (2) GitHub alone, Debian and Ubuntu together, and Debian and Ubuntu alone lead the ranking, meaning that also many CVEs have only been reported in one ecosystem without a propagation to others. The frequent propagation from Debian to Ubuntu is also expected, given the close resemblance and collaboration between these two Linux distributions. Furthermore, (3) the traceability delays, as measured by time differences between the last and first appearances of CVEs in given ecosystems, do not correlate well with the number of ecosystems present in the corresponding propagation sequences. That is, longer (shorter) sequences do not imply lengthier (faster) delays. Although reporting was omitted for brevity, it can be noted that neither the severity of the CVE-referenced vulnerabilities, as measured by the Common Vulnerability Scoring System (v. 3.1), affects the traceability delays statistically.555 https://www.first.org/cvss/v3-1/ In addition, (4) there is no clear—or at least easily interpretable—pattern in the traceability delays with respect to some particular ecosystems appearing in the sequences. Last but not least, (5) the traceability delays have been lengthy on average; the median is about two years. There are also extremely outlying traceability delays.
In light of the rather long delays, it could be perhaps contemplated whether the OSV database could (or should) be developed further into a more general platform for coordinating and otherwise handling of vulnerabilities in the open source software context [15]. Although much of development nowadays happens on GitHub, it remains unclear how well the platform works in facilitating coordination, especially when keeping in mind the large amounts of software distributed through the twenty-eight open source software ecosystems covered in the paper.

### v2

@@ 1 Introduction
The paper continues and advances the empirical vulnerability coordination research theme [10, 12] in a software ecosystem context; the interest is to better understand how vulnerabilities propagate across ecosystems. The ecosystem context places the paper into a large branch that has examined software and cyber security of particularly programming language ecosystems, including with respect to a recent increase of malware uploads to these ecosystems [1, 13]. To narrow the branch a little, the paper can be characterized to operate specifically in a cross-ecosystem context that has recently received further attention [15, 17]. In other words, the paper empirically examines the propagation of vulnerabilities across multiple open source software ecosystems. With respect to software traceability research, which too is a large branch, the paper operates in a post-release traceability context because all software observed has already been released [7, 11]. Patching and otherwise handling of vulnerabilities is a classical example of typical post-release software engineering activities [9, 19], and, hence, it is not a surprise that also vulnerabilities have been traced [2, 3, 18].
The paper’s practical relevance can be motivated by the Cyber Resilience Act (CRA) recently agreed upon in the European Union.111 Regulation (EU) 2024/2847. Among the regulation’s essential cyber security requirements are legal obligations to follow a coordinated vulnerability disclosure policy, supply security patches, and to only ship products without known vulnerabilities. As open source software is widely used also in the commercial software industry, fulfilling the obligations requires tracking, coordinating, and tracing of vulnerabilities also with respect to open source software components, including those distributed through software ecosystems. Finally and regarding the remaining structure: Section 2 presents the data and methods used, Section 3 the results, and Section 4 a conclusion and a discussion.
@@ 2.1 Data
A note should be also made about the ecosystem term; it refers to the corresponding concept used in the OSV database.333 https://google.github.io/osv.dev/data/#covered-ecosystems These ecosystems cover programming language ecosystems, such as PyPI for Python or Packagist for PHP, Linux distributions, such as Red Hat and Ubuntu, software testing frameworks, such as OSS-Fuzz, and large hosting services, such as GitHub for which GitHub Actions were merged with the main GitHub ecosystem entries.444 Although OSV uses a term Git, the term GitHub is used for clarity. However, it must be also acknowledged that the terminology used is a limitation because it is not entirely clear to which the Git entries specifically refer in all cases. This point serves also as a note that further research is needed to validate the OSV database. This terminological choice can be justified on the grounds that there is no consensus in the literature about definitions [6]. The choice also maintains coherence with OSV.
@@ 2.2 Methods
$\textmd{GitHub}\preceq\textmd{Debian}\preceq\textmd{Ubuntu},$ (4)
@@ 3 Results
The sequences elaborated in (3) and (4) provide a good way to start the dissemination of the empirical results. Thus, the top-30 sequences (in terms of frequency) are shown in Fig. 1. (For presentation purposes, the figure shows also cases with $|\sigma_{c}|=1$.) In total, a little over four thousand unique sequences were identified for the about $85$ thousand unique CVEs. This amount alone testifies that the propagation can be quite complex among the $28$ ecosystems sampled. Both the mean and median are six ecosystems per a CVE on average. However, GitHub alone leads the ranking in Fig. 1; nearly $14\%$ of the CVEs observed were reported only in GitHub and nowhere else. GitHub is also present in many traces with a $|\sigma_{c}|=2$, meaning that many other ecosystems, including both programming language ecosystems and Linux distributions, have tended to pick CVE-referenced vulnerabilities that were initially reported on GitHub.
@@
The second place in the ranking is taken by CVEs first reported in Debian and then in Ubuntu. This observation is logical because most of Ubuntu’s packages are distributed also in Debian, and largely also maintained therein. Having said that, there are also quite a few CVEs that were reported either in Debian or Ubuntu alone. Although further validation would be required, this observation could be taken to support an argument that there might be still room for better coordination and synchronization. The same point extends toward other Linux distributions who too seem to often pick CVEs that were first reported in Debian. In this regard, it also worth emphasizing that Debian, Ubuntu, Red Hat, and SUSE, among a few others, such as the Python Software Foundation who maintains PyPI, are official CVE numbering authorities (CNAs) unlike Alpine Linux, Mageia, Rocky Linux, and some others.555 https://www.cve.org/ProgramOrganization/CNAs Because the dataset’s scope was restricted to CVE-referenced vulnerabilities, and because CVEs require coordination of their own [12], the propagation sequences observed may be partially—but unlikely fully—explained by the CNAs present in the sequences.
@@
Regarding $\textmd{H}_{2}$, the Bonferroni-corrected and variance-adjusted $t$-tests indicate that for ten ecosystems out of the differences between the means are not different from zero, meaning that a presence or an absence of a given $a$ does not affect the averages of the given traceability delays.666 CRAN was excluded because only a single CVE has been reported for it. Among these ten ecosystems are GitHub, Maven, openSUSE, Packagist, PyPI, RubyGems, and Ubuntu. In contrast, statistically significant differences are present for Debian, Mageia, Red Hat, and SUSE for which their presences indicate longer delays. If the propagation sequences have involved Go or npm, the delays have been shorter. In general, these test results indicate that there is no clear-cut pattern among the ecosystems with respect to the traceability delays, including regarding the CNAs.
@@ 4 Conclusion
Regarding the results, (1) the propagation has generally been rather complex, as demonstrated by over four thousand unique propagation sequences for the $28$ ecosystems and about $85$ thousand unique CVEs observed. In terms of a frequency ranking, however, (2) GitHub alone, Debian and Ubuntu together, and Debian and Ubuntu alone lead the ranking, meaning that also many CVEs have only been reported in one ecosystem without a propagation to others. The frequent propagation from Debian to Ubuntu is also expected, given the close resemblance and collaboration between these two Linux distributions. Furthermore, (3) the traceability delays, as measured by time differences between the last and first appearances of CVEs in given ecosystems, do not correlate well with the number of ecosystems present in the corresponding propagation sequences. That is, longer (shorter) sequences do not imply lengthier (faster) delays. Although reporting was omitted for brevity, it can be noted that neither the severity of the CVE-referenced vulnerabilities, as measured by the Common Vulnerability Scoring System (v. 3.1), affects the traceability delays statistically.777 https://www.first.org/cvss/v3-1/ In addition, (4) there is no clear—or at least easily interpretable—pattern in the traceability delays with respect to some particular ecosystems appearing in the sequences. Last but not least, (5) the traceability delays have been lengthy on average; the median is about two years. There are also extremely outlying traceability delays.
Regarding future research, a good and relevant research topic would involving validating the OSV database. In other words, (a) it remains unclear how accurate and generally robust the database is. This uncertainty affects also the results reported. In addition, (b) a notable limitation of the paper is that not all fields were included in the data extraction. The affected field is worth mentioning explicitly in this regard. Likewise, (c) the focus on CVE-referenced vulnerabilities can be mentioned as a limitation, although it remains unclear whether and how cross-ecosystem empirical analysis could be pursued without unique identifiers. Given the recent surge of malware uploads to some of the ecosystems, as also archived in the OSV database [13], (d) it might make also sense to extend an empirical analysis beyond vulnerabilities. In a similar vein, (e) it would seem reasonable to extend the propagation concept. In this regard, the OSV database archives also security advisories with the references field. Taking these and possibly other traces into account would perhaps give a more complete picture of the propagation dynamics across ecosystems and open source software projects.
Addressing these topics would help at eventually tackling the enduring challenges related to data quality and vulnerability provenance meta-data [5]. The OSV database has already facilitated a better collection of meta-data but further work is required to put the meta-data collected into use. By retrieving and validating data referenced in the OSV’s traceability links, it would possible to continue toward a more elaborate propagation analysis. In light of the rather long propagation delays, it could be perhaps even contemplated whether the OSV database could (or should) be developed further into a more general platform for coordinating and otherwise handling of vulnerabilities in the open source software context [19]. Although much of development nowadays happens on GitHub, it remains unclear how well the platform works in facilitating coordination, especially when keeping in mind the large amounts of software distributed through the twenty-eight open source software ecosystems covered in the paper.